Government G levies an income tax with the following rate structure:

Percentage Rate Bracket
6 % Income from –0– to $30,000
10 Income from $30,001 to $70,000
20 Income from $70,001 to $200,000
28 Income in excess of $200,000


Taxpayer O earns $58,500 annually during years 1 through 10. Taxpayer P earns $21,500 annually during years 1 through 5 and $95,500 annually during years 6 through 10. Assume the tax rate bracket has not changed.

  1. How much total income does each taxpayer earn over the 10-year period?
  2. Compute each taxpayer’s average tax rate for the 10-year period.
  • Required A

How much total income does each taxpayer earn over the 10-year period?

  • Required B

Compute each taxpayer’s average tax rate for the 10-year period. (Round your answers to 1 decimal place.)

Average Tax Rate
Taxpayer O %
Taxpayer P %
Total Income
Taxpayer O
Taxpayer P
Answer #1

Taxpayer O earns $58,500 annually during years 1 through 10.

Total income earned by taxpayer O = 58,500 x 10

= $585,000

Annual tax paid by taxpayer O = 30,000 x 6% + (58,500 - 30,000) x 10%

= 1,800 + 2,850

= $4,650

Average tax rate for taxpayer O = Annual tax paid/Annual income

= 4,650/58,500

= 7.95%

Taxpayer P earns $21,500 annually during years 1 through 5 and $95,500 annually during years 6 through 10

Total income of taxpayer P = 21,500 x 5 + 95,500 x 5

= 107,500 + 477,500

= $585,000

Annual tax paid by taxpayer P for the years 1 to 5 = 21,500 x 6%

= $1,290

Hence, total tax paid for years 1 to 5 = 1,290 x 5

= $6,450

Annual tax paid by taxpayer P for the years 6 to 10 = 30,000 x 6% + 40,000 x 10% + (95,500 - 70,000) x 20%

= 1,800 + 4,000 + 5,100

= $10,900

Hence, total tax paid for years 6 to 10 = 10,900 x 5

= $54,500

Total tax paid by taxpayer P for 10 years = 6,450 + 54,500

= $60,950

Average tax rate for taxpayer P = Total tax paid/Total income

= 60,950/585,000

= 10.42%

Average tax rate
Taxpayer O 7.95 %
Taxpayer P 10.42 %
Total income
Taxpayer O $585,000
Taxpayer P $585,000